### v1

# On real forms of Fermat hypersurfaces

Abstract.
In this paper, we compute the number of real forms of Fermat hypersurfaces for degree $d\geq 3$ except the degree $4$ surface case, and give explicit descriptions of them.
1. Introduction
A real form of a complex algebraic variety $X$ is a real algebraic variety whose complexification is isomorphic to $X$. Now, some examples of smooth complex varieties with infinitely many nonisomorphic real forms are known. The first example was obtained by Lesieutre in [Les18] for varieties of dimension $6$, and the existence of such varieties for any dimension $n\geq 2$ was shown in [DO19]. Also, for dimension $1$, it is known that there are no such varieties. In addition, the construction of smooth complex rational affine surface with uncountably many nonisomorphic real forms, and some other results about finiteness problem of real forms is summarized in [Bot21].
Considering the real forms of specific varieties is also an interesting question. Fermat hypersurface of dimension $n$ and degree $d$ is a projective variety defined as follows:
$F^{n}_{d(,k)}:=\left(X_{0}^{d}+\cdots+X_{n}^{d}+X_{n+1}^{d}=0\right)\subset\mathbb{P}_{k}^{n+1}$
where $k$ is the base field, and we don’t write it explicitly when it is clear. In [Ogu], the following question is asked:
Question 1.1.
Find all real forms of $F^{n}_{d,\mathbb{C}}$ up to $\mathbb{R}$-isomorphisms.
When $d=1$ or $(n,d)=(1,2)$, $F^{n}_{d}$ is isomorphic to a projective space, and this has exactly $1$ real form if $n$ is even, and $2$ real forms if $n$ is odd, up to isomorphisms, which is proved in [Ben16]. Also, it is proved in [Rus02] that $F^{2}_{2}$ has exactly $4$ real forms. Note that by [BS64], for degree $2$, $F^{n}_{2}$ has only finitely many real forms up to isomorphisms as the first Galois cohomology of an affine linear algebraic group is finite.
In this paper, we give the following partial answer to the above question:
Main Theorem.
Suppose $(n,d)\neq(2,4)$ and $d\geq 3$. Then, the number of real forms of $F^{n}_{d,\mathbb{C}}$ is
$\begin{cases}\left\lfloor\frac{n}{2}\right\rfloor+2&(d\equiv 1\pmod{2})\\ \frac{(n+3)(n+5)}{8}&(d\equiv 0,\ n\equiv 1\pmod{2})\\ \frac{(n+4)(n+6)}{8}&(d\equiv 0,\ n\equiv 0\pmod{2})\end{cases}$
We prove this theorem by computing $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$. In Section 3, we deal with the case $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and in Section 4, we deal with the case $(n,d)=(1,3)$. In both cases, we have already known the automorphism group $\operatorname{\operatorname{Aut}}F^{n}_{d}$, as the former case is due to Shioda [Shi87], and the latter is due to the fact that $F^{1}_{3}$ is elliptic. So we will check the condition for an automorphism of $F^{n}_{d}$ to be a cocycle, compute a conjugation of a cocycle by an automorphism, and get this result.
We also give explicit descriptions of these real forms for the curve case in Theroem 5.1, and for the surface case in Theorem 5.2 in the way we can generalize it to higher dimensions, and give a topological description of a real locus of some real forms (in Theorem 5.3, and in particular, Theorem 5.4 for the curve case). We also show that $F_{4}^{2}$ has at least $4$ non-isomorphic real forms in Remark 5.5.
2. Preliminary about real forms
First, we review the definition of a real form and related concepts.
Definition 2.1.
Let $X$ be a complex algebraic variety and $X^{\prime}$ a real algebraic variety.
  1. (1)
    The complexification of $X^{\prime}$ is the complex algebraic variety
    ${X^{\prime}}_{\mathbb{C}}:=X^{\prime}\times_{\operatorname{Spec}\mathbb{R}}\operatorname{Spec}\mathbb{C}.$
  2. (2)
    A real form of $X$ is a real algebraic variety $X_{\mathbb{R}}$ with a $\mathbb{C}$-isomorphism $(X_{\mathbb{R}})_{\mathbb{C}}\xrightarrow[]{\sim}X$.
  3. (3)
    A real structure on $X$ is an anti-involution $\rho\colon X\to X$, i.e., an involution that makes the following diagram commute:
[…]
    where the bottom arrow is the morphism induced by the complex conjugation.
  4. (4)
    Real structures $\rho$, $\rho^{\prime}$ on $X$ are equivalent if there exists a $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\rho^{\prime}=\varphi^{-1}\circ\rho\circ\varphi$.
Theorem 2.2 ([BS64]).
Let $X$ be a complex quasi-projective variety with a real structure $\rho$ on $X$. Then, there are bijections between
  • the set of $\mathbb{R}$-isomorphism classes of real forms of $X$,
  • the set of equivalent classes of real structures on $X$, and
  • $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}_{\mathbb{C}}X)$ where the nontrivial element of $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ acts on $\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ by the conjugation of $\rho$.
It is clear by definition that there exists a bijection between the second and the third set, and existence of a bijection between the first and the third set is proved in [Se02, Chapter III, 1.3].
Remark 2.3.
In the above condition,
  • $\alpha\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ is a cocycle if and only if $(\alpha\circ\rho)^{2}=\operatorname{id}_{X}$.
  • cocycles $\alpha,\beta\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ are equivalent if and only if there exists $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\beta\circ\rho=\varphi^{-1}\circ(\alpha\circ\rho)\circ\varphi$, and we denote it by $\alpha\sim\beta$.
3. The number of real forms of Fermat hypersurfaces
Before thinking about real forms of Fermat hypersurfaces, we want to think about automorphisms of them. Clearly, there are two types of automorphisms when the base field is algebraically closed and of characteristic 00. An automorphism of the first type is the one with which we can identify $\sigma\in S_{n+2}$ defined as follows:
$F^{n}_{d}\to F^{n}_{d};[X_{0}:\cdots:X_{n+1}]\mapsto[X_{\sigma(0)}:\cdots:X_{\sigma(n+1)}]$
with $\sigma$ viewed as a permutation of $\{0,\ldots,n+1\}$.
An automorphism of the second type is the one denoted by $g(i,j)$ for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1 and defined as follows:
$F^{n}_{d}\to F^{n}_{d};[X_{0}:\cdots X_{n+1}]\mapsto[X_{0}:\cdots:\xi^{j}X_{i}:\cdots:X_{n+1}]$
with $\xi=\xi_{d}$ a primitive $d$-th root of unity.
Let $G^{n}_{d}$ be the subgroup of $\operatorname{\operatorname{Aut}}F^{n}_{d}$ generated by $g(i,j)$ for all $i,j\in\mathbb{Z}$. Then, the following result is known.
Theorem 3.1 ([Shi87]).
Let $k$ be an algebraically closed field with $\operatorname{char}k=0$, and suppose $(n,d)\neq(1,3),(2,4)$ and $d\geq 3$. Then,
$\operatorname{\operatorname{Aut}}F^{n}_{d,k}=G^{n}_{d}\cdot S_{n+2}$
where normal-$\cdot$ denotes a semidirect product.
In particular, $\operatorname{\operatorname{Aut}}F^{n}_{d,\mathbb{C}}=G^{n}_{d}\cdot S_{n+2}$.
Though this theorem is proved only in the curve case in that paper, the same argument works in higher dimensions as we can show that every automorphism is linear.
From now on, we assume $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and the base field is $\mathbb{C}$ unless otherwise stated.
Let $\tilde{\rho}\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1}$ be the complex conjugation, i.e., the $\mathbb{R}$-morphism defined by $\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X_{i}$ to $X_{i}$ for all $i$ and $c\in\mathbb{C}$ to $\bar{c}$. This is a real structure of $\mathbb{P}^{n+1}$, and it induces the real structure $\rho\colon F^{n}_{d}\to F^{n}_{d}$ on $F^{n}_{d}$.
Remark 3.2.
By straightforward calculations, for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1, $\sigma\in S_{n+2}$,
  • $g(i,j)\circ\sigma=\sigma\circ g(\sigma^{-1}(i),j)$.
  • $\sigma\circ\rho=\rho\circ\sigma$.
  • $g(i,j)\circ\rho=\rho\circ g(i,d-j)=\rho\circ g(i,j)^{-1}$.
The following result is the main step to prove the main theorem.
Proposition 3.3.
  1. (1)
    $\alpha=\mu\sigma\in\operatorname{\operatorname{Aut}}F^{n}_{d}$, with $\mu=g(0,p_{0})\cdots g(n+1,p_{n+1})\in G_{d}^{n}$, $\sigma\in S_{n+2}$, is a cocycle if and only if the following two conditions hold:
    • $\sigma$ is of type $1^{l}\cdot 2^{m}$ with $l+2m=n+2$.
    • $p_{i}\equiv p_{\sigma(i)}\pmod{d}$ if $i\in\operatorname{supp}\sigma$.
    Here, $\operatorname{supp}\sigma$ is the support of a permutation $\sigma$, i.e., $\operatorname{supp}\sigma:=\left\{i|\sigma(i)\neq i\right\}$.
  2. (2)
    Cocycles $\beta=\lambda\tau$, $\gamma$ with $\lambda\in G^{n}_{d}$ and $\tau\in S_{n+2}$ are equivalent if and only if there exist $\psi\in S_{n+2}$ and $q_{j}\in\mathbb{Z}$ for $10\leq j\leq n+$ 1 with $q_{j}\equiv q_{\sigma(j)}\pmod{2}$ for all $j$ and $q_{l}\equiv 0\pmod{2}$ for all $l\notin\operatorname{supp}\tau$ such that
    $\gamma=\psi^{-1}\cdot\left(g(0,q_{0})\cdots g(n+1,q_{n+1})\beta\right)\cdot\psi.$
Proof.
(1) By definition, $\alpha$ is a cocycle if and only if $(\mu\sigma\rho)^{2}=\operatorname{id}$. Also,
$(\mu\sigma\rho)^{2}$ $=\mu\sigma\mu^{-1}\sigma$
$=g(0,{p^{\prime}}_{0})\cdots g(n+1,{p^{\prime}}_{n+1})\cdot\sigma^{2}$
where
${p^{\prime}}_{i}=\begin{cases}0&(i\notin\operatorname{supp}\sigma)\\ p_{i}-p_{\sigma(i)}&(i\in\operatorname{supp}\sigma)\end{cases}$
Therefore, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and ${p^{\prime}}_{i}\equiv 0\pmod{d}$, and the result follows.
(2) Suppose $\beta=\lambda\tau$ be a cocycle and $\varphi=\nu\psi\in\operatorname{\operatorname{Aut}}F^{n}_{d}$ with $\lambda,\nu=g(0,r_{0})\cdots g(n+1,r_{n+1})\in G^{n}_{d},\ \tau,\psi\in S_{n+2}$. Then,
$\varphi^{-1}\circ\beta\rho\circ\varphi$
$=$ $(\psi^{-1}\nu^{-1})\cdot(\lambda\tau\rho)\cdot(\nu\psi)$
$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\tau\nu^{-1})\cdot\psi\rho$
$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\cdot g(\tau(0),d-r_{0})\cdots g(\tau(n+1),d-r_{n+1}))\cdot\tau\psi\rho$
$=$ $\psi\cdot(g(0,s_{0})\cdots g(n+1,s_{n+1})\cdot\lambda\tau)\cdot\psi\rho$
where
$s_{i}=\begin{cases}-2r_{i}&(i\notin\operatorname{supp}\tau)\\ -r_{i}-r_{\tau(i)}&(i\in\operatorname{supp}\tau)\end{cases}$
As $r_{i}$ varies for all integers, the result follows. ∎
Note that, as cocycles,
$I_{r}=\xi^{m}I_{r}$
for $m\in\mathbb{Z}$ by definition, and
$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{2i}&\\ &I_{r}\end{pmatrix},\ \begin{pmatrix}I_{r}&&\\ &0&1\\ &1&0\end{pmatrix}\sim\begin{pmatrix}I_{r}&&\\ &0&\xi^{j}\\ &\xi^{j}&0\end{pmatrix}$
for $i,j\in\mathbb{Z}$ by the proposition.
By this proposition, we can describe $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
Theorem 3.4.
Suppose $d$ is odd. Then, any cocycles are equivalent to exactly one of the following forms:
[…]
In particular, $F^{n}_{d}$ has exactly $\left\lfloor\frac{n}{2}\right\rfloor+2$ real forms.
In particular, if $d\geq 3$ is odd, then $F^{1}_{d}$ ($d\neq 3$) has exactly $2$ real forms and $F^{2}_{d}$ has $3$ real forms.
Proof.
As $d$ is odd,
$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{i}&\\ &I_{r}\end{pmatrix}$
for all $i$, and the former statement follows from Proposition 3.3. Also, as any two of $H^{r}_{n}$’s are not equivalent,
$H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})=\left\{H^{r}_{n}|0\leq r\leq n+2\right\}$
where we identify cocycles with their equivalent classes, and the latter statement follows. ∎
Theorem 3.5.
Suppose $d$ is even. Then, any cocycles are equivalent to exactly one of the forms:
$K^{s,t}_{n}:=\begin{pmatrix}\xi I_{s}&&&&&&\\ &I_{t}&&&&&\\ &&0&1&&&\\ &&1&0&&&\\ &&&&\ddots&&\\ &&&&&0&1\\ &&&&&1&0\end{pmatrix}\ (s,t\in\mathbb{Z}_{\geq 0},\ s+t\leq n+2)$
Also, $F^{n}_{d}$ has exactly $\frac{(n+3)(n+5)}{8}$ real forms for $n$ odd, and $\frac{(n+4)(n+6)}{8}$ real forms for $n$ even.
In particular, if $d\geq 3$ is even, then $F^{1}_{d}$ has exactly $3$ real forms and $F^{2}_{d}$ ($d\not=4$) has $6$ real forms.
Proof.
The former statement is clear from proposition 3.3.
By the proposition, for $s,s^{\prime},t,t^{\prime}\in\mathbb{Z}$ with $s+t=s^{\prime}+t^{\prime}$, $K^{s,t}_{n}\sim K^{s^{\prime},t^{\prime}}$ if and only if there exists a permutation matrix $\sigma$ and $a_{1}$, $\ldots$, $a_{n+2}$ with $a_{1}\equiv\cdots\equiv a_{m}\pmod{2}$ such that
$K^{s,t}_{n}\cdot\left(\sigma^{-1}K^{s^{\prime},t^{\prime}}_{n}\sigma\right)^{-1}=\begin{pmatrix}\xi^{a_{1}}&&\\ &\ddots&\\ &&\xi^{a_{n+2}}\end{pmatrix}.$
This condition is equivalent to $s=s^{\prime}$ or $s=t^{\prime}$.
Therefore,
[…]
where we identify cocycles with their equivalent classes. From this, when $n=2k-1$ for $k\in\mathbb{Z}$, the number of real forms is
$\sum_{i=0}^{k}(i+1)=\frac{(k+1)(k+2)}{2}=\frac{(n+3)(n+5)}{8},$
and when $n=2k$ for $k\in\mathbb{Z}$, the number of real forms is
$\sum_{i=0}^{k+1}(i+1)=\frac{(k+2)(k+3)}{2}=\frac{(n+4)(n+6)}{8},$
so the result follows. ∎
We omit subindexes of $H^{r}_{n}$ or $K^{s,t}_{n}$ if it is clear from the context.
4. The elliptic curve case
In this section, we consider the real forms of $F^{1}_{3}=(X_{0}^{3}+X_{1}^{3}+X_{2}^{3}=0)\subset\mathbb{P}^{2}$. First, the substitution
$X=\frac{-12X_{2}}{X_{0}+X_{1}},\ Y=\frac{36(X_{0}-X_{1})}{X_{0}+X_{1}}$
gives an elliptic curve $Y^{2}=X^{3}-432$ with $j$-invariant 00, and this can be identified with $E:=\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$ where $\zeta$ is a primitive $6$-th root of unity, which we fix in this section.
Then, any automorphisms of $E$ can be expressed by
$z\mapsto\zeta^{i}z+a$
with $i\in\mathbb{Z}$ and $a$ in the fundamental domain (cf. [Sil09]). Note that the real structure $\rho\colon F^{1}_{3}\to F^{1}_{3}$ corresponds to $E\to E$ that maps $z$ to $\bar{z}$ where we identify $z\in\mathbb{C}$ with the image of $z$ by the canonical surjection $\mathbb{C}\to\mathbb{C}/(\mathbb{Z}+\zeta\mathbb{Z})$. We also denote it by $\rho\colon E\to E$.
Proposition 4.1.
We fix $\xi\in\mathbb{C}$ so that $\xi^{2}=\zeta$.
  1. (1)
    $\alpha\colon E\to E;\ z\mapsto\zeta^{i}z+a$ is a cocycle if and only if there exists $r\in\mathbb{R}$ such that $a=\sqrt{-1}\xi^{i}r$.
  2. (2)
    Any cocycles are equivalent to $z\mapsto z$ or $z\mapsto-z$, which are not equivalent to each other.
Proof.
(1) By definition, $\alpha$ is a cocycle if and only if $(\alpha\circ\rho)^{2}=\operatorname{id}$.
$(\alpha\circ\rho)^{2}(z)$ $=\zeta^{i}(\overline{\zeta^{i}\bar{z}+a})+a$
$=z+(\zeta^{i}\bar{a}+a)$
Therefore, the result follows.
(2) Let $\sigma\colon E\to E;\ z\mapsto\zeta^{i}z+a$ be a cocycle and $\tau\colon E\to E;\ z\mapsto\zeta^{j}z+b$. Then, $\tau^{-1}\colon z\mapsto\zeta^{-j}z-\zeta^{j}b$. Therefore,
$(\tau^{-1}\circ(\sigma\rho)\circ\tau)(z)$
$=$ $\tau^{-1}(\zeta^{i}(\overline{\zeta^{j}z+b})+a)$
$=$ $\zeta^{i-2j}\bar{z}+\zeta^{-j}a-\zeta^{j}b+\zeta^{i-j}\bar{b}$
First, $z\mapsto z$ and $z\mapsto-z$ are not equivalent to each other by this calculation.
Also, in the above, we can write $a=\sqrt{-1}\xi^{i}x$ for $x\in\mathbb{R}$ by (1). Then, by setting
$j=0,\ b=\frac{1+\sqrt{-1}}{2}\xi^{i}x,$
we see that $z\to\zeta^{i}z+a$ is equivalent to $z\to\zeta^{i}z$, so the result follows. ∎
Note that the negation map in the above corresponds to a linear automorphism
$F^{1}_{3}\to F^{1}_{3};\ [X_{0}:X_{1}:X_{2}]\mapsto[X_{1}:X_{0}:X_{2}]$
by taking $[1:-1:0]$ as a unit. From this, we immediately get the following result:
Theorem 4.2.
In $F^{1}_{3}$, any cocycles are equivalent to $H^{3}_{1}$ or $H^{1}_{1}$.
In particular, $F^{1}_{3}$ has exactly $2$ real forms.
5. Explicit description of real forms of Fermat hypersurfaces
First, we consider the curve case $F^{1}_{d}=(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$.
Theorem 5.1.
Assume that $d\geq 3$. Then, the correspondence between cocycles and real forms of $F^{1}_{d}$ is given by the following table.
cocycle real form
$K^{0,3}_{1}$ (or $H^{0}_{1}$) $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}=0)\subset\mathbb{P}^{2}_{\mathbb{R}}$
$K^{1,2}_{1}$ $(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}=0)\subset\mathbb{P}^{2}_{\mathbb{R}}$
$K^{0,1}_{1}$ (or $H^{2}_{1}$) $(Y_{0}^{d}+(Y_{1}+\sqrt{-1}Y_{2})^{d}+(Y_{1}-\sqrt{-1}Y_{2})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}$
Proof.
First, the real form corresponding to $K^{0,3}$ is clear.
Next, we consider the real form corresponding to $K^{1,2}$. The invariant ring of the homomorphism $\alpha\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $\xi_{d}X_{0}$, $X_{1}$ to $X_{1}$, and $X_{2}$ to $X_{2}$ is $\mathbb{R}[\xi_{2d}X_{0},X_{1},X_{2}]$ with $\xi_{2d}^{2}=\xi_{d}$.
In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\mu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
$f(X_{0},X_{1},X_{2})=\mu(f)(\xi_{2d}X_{0},X_{1},X_{2}),$
and $\alpha(f)=f$ if and only if $\mu(f)$ is a real coefficient polynomial, as $\xi_{2d}X_{0}$, $X_{1}$, $X_{2}$ are invariant under $\alpha$.
Therefore, by setting
$Y_{0}:=\xi_{2d}X_{0},\ Y_{1}:=X_{1},\ Y_{2}:=X_{2},$
we see that the corresponding real form is
$(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
Finally, we consider the real form corresponding to $K^{1,2}$. The invariant ring of the homomorphism $\beta\colon\mathbb{C}[X_{0},X_{1},X_{2}]\to\mathbb{C}[X_{0},X_{1},X_{2}]$ that sends $c\in\mathbb{C}$ to $\bar{c}$, $X_{0}$ to $X_{0}$, $X_{1}$ to $X_{2}$ and $X_{2}$ to $X_{2}$ is $\mathbb{R}[X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})]$.
We can show this by a similar argument as above. In fact, for all $f(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$, there exists a unique polynomial $\nu(f)(X_{0},X_{1},X_{2})\in\mathbb{C}[X_{0},X_{1},X_{2}]$ such that
$f(X_{0},X_{1},X_{2})=\nu(f)(X_{0},X_{1}+X_{2},(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2})),$
and $\beta(f)=f$ if and only if $\nu(f)$ is a real coefficient polynomial.
Therefore, by setting
$Y_{0}:=X_{0},\ {Y^{\prime}}_{1}:=X_{1}+X_{2},\ {Y^{\prime}}_{2}:=(1-\sqrt{-1})(X_{1}-\sqrt{-1}X_{2}),$
and also, setting
$Y_{1}:=\frac{{Y^{\prime}}_{1}}{2},\ Y_{2}:=\frac{{Y^{\prime}}_{2}-{Y^{\prime}}_{1}}{2},$
we see that the corresponding real form is
$(Y_{0}^{d}+(Y_{1}+\sqrt{-1})^{d}+(Y_{1}-\sqrt{-1})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{2}.$
In the surface case, by using the same argument as above, we get the following result:
Theorem 5.2.
Assume $d\geq 3,\neq 4$. Then, the correspondence between cocycles and real forms of $F^{2}_{d}$ is given by the following table.
cocycle real form
$K^{0,4}_{2}$ (or $H^{4}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{1,3}_{2}$ $(-Y_{0}^{d}+Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{2,2}_{2}$ $(-Y_{0}^{d}-Y_{1}^{d}+Y_{2}^{d}+Y_{3}^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{0,2}_{2}$ (or $H^{2}_{2}$) $(Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{1,1}_{2}$ $(-Y_{0}^{d}+Y_{1}^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
$K^{0,0}_{2}$ (or $H^{0}_{4}$) $((Y_{0}+\sqrt{-1}Y_{1})^{d}+(Y_{0}-\sqrt{-1}Y_{1})^{d}+(Y_{2}+\sqrt{-1}Y_{3})^{d}+(Y_{2}-\sqrt{-1}Y_{3})^{d}=0)\subset\mathbb{P}_{\mathbb{R}}^{3}$
In the same way, we can extend it to the general case. In particular, all real forms of $F_{d}^{n}$ are the $n$-dimensional hypersurface of degree $d$.
Next, we will give a topological description of a real locus of these real forms.
Theorem 5.3.
  1. (1)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{n+1}^{d}=0)$
    with $d$ even is $\emptyset$.
  2. (2)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{n}^{d}-X_{n+1}^{d}=0)$
    with $d$ even is homeomorphic to $S^{n}$.
  3. (3)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{i}^{d}-X_{i+1}^{d}-\cdots-X_{n+1}^{d}=0)$
    with $d$ even is homeomorphic to $\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$.
  4. (4)
    The real locus of
    $\left(\sum_{k=0}^{i-1}X_{k}^{d}+\sum_{l=0}^{\frac{n-i}{2}}\left(\left(X_{i+2l}+\sqrt{-1}X_{i+2l+1}\right)^{d}+\left(X_{i+2l}-\sqrt{-1}X_{i+2l+1}\right)^{d}\right)=0\right)$
    with $d$ odd and $i>0$ is homeomorphic to $\mathbb{P}^{n}(\mathbb{R})$.
  5. (5)
    The real locus of
    $(X_{0}^{d}+\cdots+X_{n-1}^{d}+(X_{n}+\sqrt{-1}X_{n+1})^{d}+(X_{n}-\sqrt{-1}X_{n+1})^{d}=0)$
    with $d$ even is homeomorphic to $\underbrace{S^{n}\sqcup\cdots\sqcup S^{n}}_{\frac{d}{2}}$.
Proof.
In each case, let $Z$ be the real locus.
(1) is clear.
(2) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $x_{n+1}\neq 0$. Therefore, we can define a continuous map $Z\to S^{n}$ as follows:
$[x_{0}:\cdots:x_{n}:1]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots+x_{n}^{2}}}(x_{0},\ldots,x_{n}).$
As it is a bijection, it is also a homeomorphism, and thus, $Z\approx S^{n}$.
(3) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{0},\ldots,x_{i})\neq(0,\ldots,0)$, $(x_{i+1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Thus, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto([x_{0}:\cdots:x_{i}],[x_{i+1}:\cdots:x_{n+1}]).$
As it is a bijection, it is also a homeomorphism, and thus, $Z\approx\mathbb{P}^{i}(\mathbb{R})\times\mathbb{P}^{n-i}(\mathbb{R})$.
(4) Note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{1},\ldots,x_{n+1})\neq(0,\ldots,0)$. Therefore, we can define a continuous function $Z\to\mathbb{P}^{n}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{1}:\cdots:x_{n+1}].$
As $d$ is odd, this is a bijection, and thus, a homeomorphism. So $Z\approx\mathbb{P}^{n}(\mathbb{R})$.
(5) First, note that $[x_{0}:\cdots:x_{n+1}]\in Z$ implies $(x_{n},x_{n+1})\neq(0,0)$. Therefore, we can define a continuous map $Z\to\mathbb{P}^{i}(\mathbb{R})$ as follows:
$[x_{0}:\cdots:x_{n+1}]\mapsto[x_{n}:x_{n+1}].$
Then, this image has $\frac{d}{2}$ connected components
$E_{m}:=\left\{[\cos\theta:\sin\theta]\bigg{|}\frac{(4m-1)\pi}{2d}\leq\theta\leq\frac{(4m+1)\pi}{2d}\right\}$
for $10\leq m\leq\frac{d}{2}-$ 1.
In fact, $[x_{n}:x_{n+1}]\in f(Z)$ if and only if $(x_{n}+\sqrt{-1}x_{n+1})^{d}+(x_{n}-\sqrt{-1}x_{n+1})^{d}\leq 0$, and
$(\cos\theta+\sqrt{-1}\sin\theta)^{d}+(\cos\theta-\sqrt{-1}\sin\theta)^{d}=2\cos d\theta.$
Also, for each $m$, we can define a continuous map $f^{-1}(E_{m})\to S^{n}$ defined as follows:
$[x_{0}:\cdots:x_{n-1}:\cos\theta:\sin\theta]\mapsto\frac{1}{\sqrt{x_{0}^{2}+\cdots x_{n-1}^{2}+\sin d\theta}}(x_{0},\ldots,x_{n-1},\sin d\theta)$
Note that this is well-defined as $d$ is even. As this is a bijection, it is also a homeomorphism, and thus $Z\approx\underbrace{S^{n}\sqcup\cdots\sqcup S^{n}}_{\frac{d}{2}}$. ∎
In particular, we determined the real locus of real forms for the curve cases:
Theorem 5.4.
  1. (1)
    Suppose $d$ is odd. Then, the real loci of
    $(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0),\ (X_{0}^{d}+(X_{1}+\sqrt{-1}X_{2})^{d}+(X_{1}-\sqrt{-1}X_{2})^{d}=0)$
    are both homeomorphic to $\mathbb{P}^{n}(\mathbb{R})$.
  2. (2)
    Suppose $d$ is even. Then,
    • the real locus of $(X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$ is $\emptyset$.
    • the real locus of $(-X_{0}^{d}+X_{1}^{d}+X_{2}^{d}=0)$ is homeomorphic to $S^{1}$.
    • the real locus of $(X_{0}^{d}+(X_{1}+\sqrt{-1}X_{2})^{d}+(X_{1}-\sqrt{-1}X_{2})^{d}=0)$ is homeomorphic to $\underbrace{S^{1}\sqcup\cdots\sqcup S^{1}}_{\frac{d}{2}}$.
Remark 5.5.
By considering real loci, we see that $F^{2}_{4}$ has at least $4$ non-isomorphic real forms
  • $(X_{0}^{4}+X_{1}^{4}+X_{2}^{4}+X_{3}^{4}=0)$,
  • $(-X_{0}^{4}+X_{1}^{4}+X_{2}^{4}+X_{3}^{4}=0)$,
  • $(-X_{0}^{4}-X_{1}^{4}+X_{2}^{4}+X_{3}^{4}=0)$, and
  • $(X_{0}^{4}+X_{1}^{4}+(X_{2}+\sqrt{-1}X_{3})^{4}+(X_{2}-\sqrt{-1}X_{3})^{4}=0)$.
Note that $F^{2}_{4}$ is a K3 surface and thus has only finitely many real forms up to isomorphisms by [CF20]. However, unlike the other cases $d\geq 3$, the automorphism group of $F^{2}_{4}$ is still unknown, and it is known to be a very difficult problem to determine the automorphism group of $F^{2}_{4}$. For this reason, it is also a very challenging problem to determine all real forms of $F^{2}_{4}$ or to see at least if there is a real form of $F^{2}_{4}$ other than these four or not.
Remark 5.6.
Let $Q^{r,s}$ be a hypersurface defined as follows:
$Q^{r,s}:=(X_{0}^{2}+\cdots+X_{r-1}^{2}-X_{r}^{2}-\cdots-X_{r+s-1}^{2}=0)\subset\mathbb{P}^{r+s-1}.$
Then, $Q^{r,s}$ with $r+s=n+2$ is a real form of $F^{n}_{2}$. Also, by considering real loci, we see that $Q^{r,s}$ and $Q^{t,u}$ with $r+s=t+u=n+2$ are $\mathbb{R}$-isomorphic if and only if $r=t$ or $r=u$.
However, there exists a real form of $F^{n}_{2}$ other than these forms. For example, by [Rus02], any real forms of $F^{2}_{2}$ are exactly $\mathbb{R}$-isomorphic to $Q^{4,0}$, $Q^{3,1}$, $Q^{2.2}$, or $Q^{3,0}\times Q^{2,1}$.